OpenStudy (anonymous):

Which of the following is greatest? thirteen twenty-fifths zero point four five forty-three one hundredths twenty-nine fiftieths

OpenStudy (anonymous):

I'd put each in the same form...all fractions or all decimals. Doing this will let you know which one is the greatest

OpenStudy (anonymous):

OK any other suggestions.

OpenStudy (anonymous):

13/25=0.52 0.45 0.43 29/50=0.58 when converted to all decimals, it's easy to see

OpenStudy (anonymous):

Awh I see now that was quite simple

OpenStudy (anonymous):

13/25=52/100 45/100 43/100 58/100 Here I used a common denominator so I could compare which is the greatest.